OpenStudy (yacoub1993):

Find the specified vector or scalar. u = -4i + 1j and v = 4i + 1j; Find ‖u+v‖. A. √34 B. 8 C. 2 D. 5

OpenStudy (yacoub1993):

@amistre64 Help Please

OpenStudy (amistre64):

les start by adding the vectors; combine like terms

OpenStudy (yacoub1993):

i have no idea how to do these kind of math Can you help me out with it

OpenStudy (amistre64):

u = -4i + 1j and v = 4i + 1j u+v = -4i + 1j + 4i + 1j lets start .... by combining like terms

OpenStudy (yacoub1993):

u+v = -4i +4i + 2j

OpenStudy (yacoub1993):

is it like this

OpenStudy (amistre64):

thats part of it ... now combine the i parts

OpenStudy (yacoub1993):

if i add -4j to 4 it will be 0j

OpenStudy (amistre64):

alittle bit of typos there ... it might help to consider that ij is a generic notation and can simply be restated as xy u+v = -4x +4x + 2y this is a suitable rendition; u+v = 0x + 2y

OpenStudy (amistre64):

now the question is asking: how long is this vector? or rather, how far is (0,2) from the origin?

OpenStudy (yacoub1993):

ohh i get it

OpenStudy (amistre64):

it is pretty basic once you know the notation :)